Reports 1-1 of 1 Clear search Modify search
AdV-COM (AdV commissioning (1st part) )
ruggi - 13:24 Saturday 05 October 2019 (47104) Print this report
BS vertical position

A scan of BS vertical position has been performed, while a line at 26 Hz was injected on BS TX, and it was visible in DARM. It has been found that BS_TX --> DARM coupling can be reduced if BS is rised a lot. The scan has been stopped after a shift of 5 mm, but the minimun coupling was not yet reached. The projection of the zero crossing says that 2.7 mm were still missing (fig 1).

We use to consider the angular to longitudinal coupling as a good measurement of the beam centering on the mirror, so it would be better to change BS vertical position in order to have the mirror correctly centered. Yesterday BS has been put back almost in the previous position, in order not to change the tuning of the vertical attenuation stages. The correct way to do the shift is by the top stage screw (chiocciolone). It is worth to remark that no relevant changes has been observed in the sensitivity during the scan: the only thing which clearly changes is the coupling of BS TX angular noise, but it is currently negligible.

The optimal BS vertical position is a bit too high, for the constrains determined by F7 sensing/actuation. But also NI and WI setting are a bit high with respect to F7 sensing, so there would be the possibility to align everything in a different position of the input beam (let's say: 2 or 3 mm lower). The impact on DET has to be checked.

Last remark: also PR is working in a very low position with respect to F7 sensors. For PR, a simple scan of the vertical position is impossible, because any small shift requires an adjustment of the input beam inclination. This means that currently we do not have a clear confirmation of the centering correctness of PR, and an activity devoted to verify/optimize that is a bit long and invasive (one entire shift, at least). But I think October stop is the good time to do it.

 

Images attached to this report
Comments to this report:
ruggi - 15:33 Tuesday 08 October 2019 (47149) Print this report

Since yesterday BS is working 4 mm above the usual position. The adjustment has been done partially by the top screw, partially adjusting the set point of some vertical attenuation stage, which was not working in the optimal position. The impact of those adjustments on the susp performance is supposed to be null (and in fact, nothing special has been observed). The impact on ITF performance should be the one already said: a lowering of BS_TX angular noise coupling (which was already negligible).

In order to put BS in the position of minimum angular noise coupling, 3 mm are still missing.

Images attached to this comment
Search Help
×

Warning

×